OpenStudy (anonymous):

Find the slope of the line passing through the points (0, –5) and (4, 2).

OpenStudy (anonymous):

This should be in math section. Anyway use; \[(y2-y1)/(x2-x1)\] This should be your slope.

OpenStudy (rina.r):

X2 Y2, X1 Y1 0 -5 , 4 2 then (0-(-5))/ (4-2) = 5 / 2 =2.5